the drawing, the weight of the block on the table is 422 N and that of the hanging block is 185 N. Ignoring all 

frictional effects and assuming the pulley to be massless, 


1
Expert's answer
2021-06-13T07:08:51-0400

So, assume that


In the drawing, the weight of the block on the table is 422 N and that of the hanging block is 185 N. Ignoring all frictional effects and assuming the pulley to be massless, find (a) the acceleration of the two blocks and (b) the tension in the cord.






"m_1=422/9.81=43\ (kg)"

"m_2=185/9.81=18.9\ (kg)"


(a) "m_1a=T" and "m_2g-T=m_2a\to 185-T=m_2a"


"185-m_1a=m_2a\to a=\frac{185}{m_1+m_2}=\frac{185}{43+18.9}=2.99\ (m/s^2)"


(b) "T=m_1a=43\cdot2.99=128.5\ (N)"